Assume the GDP is $6,000, private domestic savings are $1,100, and the government budget deficit is $200. Consumption is $3,800, and the trade deficit is $100.

Find: (a) Investments (I), (b) Government spending (G)
Answer #1

Answer

Private saving=Y-T-C

1100=6000-T-3800

T=$1100


Public saving=(T-G)=budget surplus-budget deficit

-200=T-G

-200=1100-G

G=1300

The government spending is $1300
==========

b)

Y=C+I+G+NX

NX=net export =trade surplus =- trade defict =-100

6000=3800+I+1300-100

I=6000-5000=1000

the investment is $1000